# Neutrino Heating in 1D, 2D, and 3D core-collapse supernovae: characterizing the explosion of high-compactness stars

### v1

@@ 6 Summary and Conclusions
From simple fits (whose details can be found in Appendix B) we were able to derive a curve describing how $\dot{Q}_{\nu}^{\rm max}$ depends on the pre-supernova compactness, and showed that multi-dimensional simulations and 1D+ models closely follow said curve. The rapid increase of $\dot{Q}_{\nu}^{\rm max}$ as a function of compactness shows that high-compactness progenitors are able to generate very strong neutrino heating, and therefore despite the larger mass accretion rates draining energy from the shock, the shock is always successfully revived. As shown by recent multi-dimensional simulations (Powell et al., 2021; Chan et al., 2018; Burrows et al., 2023; Sykes & Müller, 2024; Eggenberger Andersen et al., 2024; Burrows et al., 2024a) after shock revival, even if the formation of a black hole is very likely, shock breakout is quite likely, and the ensuing explosion can have energies ranging from 0.01-3.5 $10^{51}$ erg/s.
We also analyzed the interplay between three different timescales of the problem: $t_{\rm max}$, i.e. the time when $\dot{Q}_{\nu}^{\rm max}$ is reached; $t_{\rm Si/O}$, i.e. the time when the Si/O interface is accreted through the shock; $t_{\rm expl}$, i.e. the time when the shock crossed 500 km. We showed that, except for a few special cases, $t_{\rm max}\approx t_{\rm Si/O}$. This shows that after the accretion of the Si/O interface neutrino heating significantly drops, and therefore achieving an explosion after that becomes much harder. Therefore, this explains why the accretion of the Si/O interface can be so important in determining the explosion (Boccioli et al., 2023; Wang et al., 2022). However, we also showed that the Si/O interface only influences the explosion for progenitors with $\xi_{2.0}<0.5$ whereas it has a much smaller impact on high-compactness ones with $\xi_{2.0}>0.5$, as also recognized by Boccioli & Fragione (2024) and Boccioli (2024b).
@@
With the advent of important missions such as the Vera C. Rubin Observatory, Cosmic Explorer, and Einstein Telescope, it will become feasible to access observational signatures of high-compactness progenitors, and therefore a better understanding of their explosion is needed. In this paper, we have begun to analyze the early explosion phase and explained why the early shock revival is always successful for high-compactness progenitors. We also showed that, in the future, 1D+ models can be used to carry out similar studies of large sets of simulations, given their agreement with neutrino heating in multi-D models, particularly for high-compactness progenitors. Notice that, however, after shock revival, the continued accretion and the asymmetry in explosion and neutrino emission introduce viewing angle effects and mixing not captured otherwise (Sandoval et al., 2021; Wongwathanarat et al., 2015; Vartanyan et al., 2024; Müller et al., 2018; Stockinger et al., 2020). Regarding high-compactness progenitors, there are still some differences among simulations from different groups. Some obtain successful shock revival and late-time shock expansion, yielding successful explosions (Burrows et al., 2024a), whereas others obtain either a successful shock revival followed by a very early black-hole formation that stops the explosion, or no shock revival at all (Kuroda et al., 2018; Walk et al., 2020; Summa et al., 2018). This further shows that more work on the later phases of the explosion is needed, and the first efforts have already begun (Sykes & Müller, 2024; Eggenberger Andersen et al., 2024; Burrows et al., 2024a; Stockinger et al., 2020).
@@ Acknowledgements
The authors would like to thank Hans-Thomas Janka and Bernhard Müller for their useful comments. L.B. would like to thank Jeremiah Murphy and Mariam Gogilashvili for fruitful discussions. The authors thank Xingzao Li and Oliver Eggenberger Andersen who have simulated many of the 2D FLASH simulations in this work. L.B. is supported by the U.S. Department of Energy under Grant No. DE-SC0004658. L.B. would like to thank the N3AS center for their hospitality and support. E.O. is supported by the Swedish Research Council (Project No. 2020-00452). Some of the computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
@@ Appendix A Neutrino heating in the gain region
which holds for electron neutrinos and antineutrinos, where the integral is evaluated in the gain region, $m_{u}=1.66\times 10^{-24}$ g is the atomic mass unit, and $\langle\sigma_{\rm abs}\rangle$ is the average cross section for neutrino absorption in the gain region. Notice that we assumed that the neutrino spectrum is forward peak, so that the integral over all neutrino angles is 1. The cross section for neutrino absorption, which is to a good approximation the same for neutrinos and antineutrinos, is defined in Eq. 19 of (Janka, 2001) as:
@@
This is motivated by the fact that progenitors with larger compactnesses have larger densities in the gain region, and therefore neutrino heating extends to larger radii, hence one expects $R^{*}$ to be slightly larger for higher-compactness progenitors. In the original work from Janka (2012), $R^{*}$ was taken to be simply the gain radius $R_{\rm g}$. To derive Eq. (14) we considered that the net neutrino heating $\dot{Q}_{\nu}$ is to a good approximation some fraction of the total heating $\dot{Q}_{\nu}^{+}$, and in this work we consider that fraction to be 0.6.
@@ Appendix B Dependence of the maximum of neutrino heating on compactness
The largest difference with the 1D simulations is however that $t_{\rm max}$ varies for different progenitors. Specifically, as shown in Section 4.1, $t_{\rm max}$ is roughly equal to the time $t_{\rm Si/O}$ when the Si/O interface is accreted through the shock. This is true up to compactnesses of $\sim\xi_{2.0}0.72$. For progenitors with larger compactnesses than that, $t_{\rm max}$ is reached at $\sim 300$ ms after bounce, i.e. the end of the accretion phase which corresponds to when the shock is revived and mass accretion onto the PNS shuts off.

### v2

@@ 6 Summary and Conclusions
From simple fits (whose details can be found in Appendix B) we were able to derive a curve describing how $\dot{Q}_{\nu}^{\rm max}$ depends on the pre-supernova compactness, and showed that multi-dimensional simulations and 1D+ models closely follow said curve. The rapid increase of $\dot{Q}_{\nu}^{\rm max}$ as a function of compactness shows that high-compactness progenitors are able to generate very strong neutrino heating, and therefore despite the larger mass accretion rates draining energy from the shock, the shock is always successfully revived. As shown by recent multi-dimensional simulations (Chan et al., 2018; Powell et al., 2021; Burrows et al., 2023; Sykes & Müller, 2025; Eggenberger Andersen et al., 2025; Burrows et al., 2024a) after shock revival, even if the formation of a black hole is very likely, shock breakout is quite likely, and the ensuing explosion can have energies ranging from 0.01-3.5 $10^{51}$ erg/s.
We also analyzed the interplay between three different timescales of the problem: $t_{\rm max}$, i.e. the time when $\dot{Q}_{\nu}^{\rm max}$ is reached; $t_{\rm Si/O}$, i.e. the time when the Si/O interface is accreted through the shock; $t_{\rm expl}$, i.e. the time when the shock crosses 500 km. We showed that, except for a few special cases, $t_{\rm max}\approx t_{\rm Si/O}$. This shows that after the accretion of the Si/O interface neutrino heating significantly drops, and therefore achieving an explosion after that becomes much harder. Therefore, this explains why the accretion of the Si/O interface can be so important in determining the explosion (Boccioli et al., 2023; Wang et al., 2022). However, we also showed that the Si/O interface only influences the explosion for progenitors with $\xi_{2.0}<0.5$ whereas it has a much smaller impact on high-compactness ones with $\xi_{2.0}>0.5$, as also recognized by Boccioli & Fragione (2024) and Boccioli et al. (2025).
@@
With the advent of important missions such as the Vera C. Rubin Observatory, Cosmic Explorer, and Einstein Telescope, it will become feasible to access observational signatures of high-compactness progenitors, and therefore a better understanding of their explosion is needed. In this paper, we have begun to analyze the early explosion phase and explained why the early shock revival is always successful for high-compactness progenitors. We also showed that, in the future, 1D+ models can be used to carry out similar studies of large sets of simulations, given their agreement with neutrino heating in multi-D models, particularly for high-compactness progenitors. Notice that, however, after shock revival, the continued accretion and the asymmetry in explosion and neutrino emission introduce viewing angle effects and mixing not captured otherwise (Wongwathanarat et al., 2015; Müller et al., 2018; Stockinger et al., 2020; Sandoval et al., 2021; Vartanyan et al., 2025). Regarding high-compactness progenitors, there are still some differences among simulations from different groups. Some obtain successful shock revival and late-time shock expansion, yielding successful explosions (Burrows et al., 2024a), whereas others obtain either a successful shock revival followed by a very early black-hole formation that stops the explosion, or no shock revival at all (Kuroda et al., 2018; Summa et al., 2018; Walk et al., 2020). This further shows that more work on the later phases of the explosion is needed, and the first efforts have already begun (Stockinger et al., 2020; Sykes & Müller, 2025; Eggenberger Andersen et al., 2025; Burrows et al., 2024a).
@@ Acknowledgements
The authors would like to thank the anonymus referee for significantly contributing to the improvement of this manuscript. They would also like to thank Hans-Thomas Janka and Bernhard Müller for their useful comments. L.B. would like to thank Jeremiah Murphy and Mariam Gogilashvili for fruitful discussions. The authors thank Xingzao Li and Oliver Eggenberger Andersen who have simulated many of the 2D FLASH simulations in this work. L.B. is supported by the U.S. Department of Energy under Grant No. DE-SC0004658 and SciDAC grant, DE-SC0024388. L.B. would like to thank the N3AS center for their hospitality and support. E.O. is supported by the Swedish Research Council (Project No. 2020-00452). D.K. is supported in part by a grant from the Simons Foundation (622817DK). Some of the computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. This research benefited from support from the Gordon and Betty Moore Foundation through Grant GBMF5076.
@@ Appendix A Neutrino heating in the gain region
which holds for electron neutrinos and antineutrinos, where the integral is evaluated in the gain region, $m_{u}=1.66\times 10^{-24}$ g is the atomic mass unit, and $\langle\sigma_{\rm abs}\rangle$ is the average cross section for neutrino absorption in the gain region. Notice that we assumed that the neutrino spectrum is forward peak, so that the integral over all neutrino angles is 1. The cross section for neutrino absorption, which is to a good approximation the same for neutrinos and antineutrinos, is defined in Eq. 19 of Janka (2001) as:
@@
This is motivated by the fact that progenitors with higher compactnesses have larger densities in the gain region, and therefore neutrino heating extends to larger radii, hence one expects $R^{*}$ to be slightly larger for higher-compactness progenitors. In the original work from Janka (2012), $R^{*}$ was taken to be simply the gain radius $R_{\rm g}$. To derive Eq. (15) we considered that the net neutrino heating $\dot{Q}_{\nu}$ is to a good approximation some fraction of the total heating $\dot{Q}_{\nu}^{+}$, and in this work we consider that fraction to be 0.6.
@@ Appendix B Dependence of the maximum of neutrino heating on compactness
The largest difference with the 1D simulations is however that $t_{\rm max}$ varies for different progenitors. Specifically, as shown in Section 4.1, $t_{\rm max}$ is roughly equal to the time $t_{\rm Si/O}$ when the Si/O interface is accreted through the shock. This is true up to compactnesses of $\sim\xi_{2.0}0.72$. For progenitors with higher compactnesses than that, $t_{\rm max}$ is reached at $\sim 300$ ms after bounce, i.e. the end of the accretion phase which corresponds to when the shock is revived and mass accretion onto the PNS shuts off.